# How Far Can Transformers Reason? The Locality Barrier and Inductive Scratchpad

### v1

@@ B.4 Length generalization for the parity task
where <pointer> determines the position of the $i$th bit and <value of the $i$th bit> is the value of the $i$th bit (which can be retrieved using the pointer) and <parity of the first $i$ bits> is the cumulative parity of the first $i$ bits that can be computed using its value in the previous state (parity of the first $i-1$ bits) and the value of the $i$th bit. For the last state, state $n+1$, we use [$d_{\mathrm{emb}}$]_,<final parity><EOS> where final parity is equal to the parity of the first $n$ bits that is computed in state $n$. So, one can easily check that this scratchpad is of low locality and each state can be easily computed using the previous state and question. As an example consider input _01_10_0__1_ again. The inductive scratchpad for this example can be written as
@@ B.5 Length generalization for the addition task
For both addition methods, we sampled numbers with $n\leq N_{\mathrm{train}}$ digits with a probability proportional to $n$. Nevertheless, we did not observe much dependency on the distribution. Also, note that in the design of scratchpads for both addition and parity, we have used techniques such as inserting random spaces and embedding in a fixed dimension. We note that this is unavoidable for the input as we have used absolute positional embedding for it. Nevertheless, we expect that by using an inductive scratchpad along with relative positional embedding and pretraining, one would be able to achieve the same length generalization with fewer assumptions on the format of the input.
@@ C.1 Architecture and datasets
For the cycles task, we use $1000$ node names formatted like v123. For simplicity of analysis, we regard each node name as a single token. Other than that and for the other tasks, we treat each character as a single token.
@@ C.3 Comparison with relative positional embeddings
Note that in the analysis above we assumed a fixed number of tokens per state. If the number of tokens in each state varies, issues like (2) and (3) above can become more difficult to handle for a model that only relies on relative positional embedding. To sum up, there is a minor similarity between relative positional embeddings and the idea of induction in the scratchpad. Factors such as pretraining and the recency bias of attention may promote inductive behavior in the model to some extent. Nevertheless, there is no reason to think that the model can implement the inductive behavior relying only on relative positional embeddings for a general inductive task. In general, one can use the inductive scratchpad idea along with relative positional embedding to get better length generalization. Note that inductive scratchpad can generalize on the number of training steps. However, understanding an input with growing length still requires relying on relative positional embeddings.
@@ Appendix E Comment on Lemma 1
Therefore, even for sets of size $n$, the mutual information is exponentially low, implying that $\mathrm{loc}(D)$ is greater than $n+1$.
@@ Appendix F Discussion on circuit complexity connections
Connection between locality and $NC^{0}$. We next show that, putting aside the histogram knowledge and using constant alphabet size, having low locality means having correlations with $NC^{0}$ circuits, a class of circuit weaker than $TC^{0}$ by constraining the number of fan-in to be constant (and not allowing for threshold gates).
@@
Let $P_{X}$ be a probability distribution over $\{0,1\}^{n}$ and $f\colon\{0,1\}^{n}\rightarrow\{0,1\}$ be a function. $f$ correlates non-trivially with a function computable in $NC^{0}$ if and only if $\mathrm{Loc}(f)=O_{n}(1)$ when $X\sim P_{X}$.
@@
The $NC^{0}$ functions are exactly the binary functions that only depend on a constant number of input bits. So, for any $NC^{0}$ function $g$, there exists a set of input bits $S$ such that $|S|=O(1)$ and $g(X)$ only depends on the restriction of $X$ to $S$. That means that if $f$ has a nontrivial correlation with $g$ then $f(X)$ has a nontrivial correlation with $X[S]$ and thus has locality at most $|S|=O(1)$.
Now, assume that $\mathrm{Loc}(f)=O_{n}(1)$ and choose $S$ with $|S|=O(1)$ such that $I(X[S],f(X))=n^{-O(1)}$. Next, define the function $g^{\prime}\colon\{0,1\}^{n}\rightarrow[0,1]$ such that $g^{\prime}(x)=\mathbb{E}[f(X)|X[S]=x[S]]$ for all $x$ and $g\colon\{0,1\}^{n}\rightarrow\{0,1\}$ such that
@@ Appendix G More experiments with ChatGPT
where the answer is true. Note that to answer this question correctly one has to combine at least $n+1$ relations. Thus, the locality of the task is always larger than $n$. (The exact locality would depend on the tokenization.) We found out that ChatGPT (GPT3.5) fails at this task even for $n=1$ (simplest case). Note that when working with the GPT3.5 model we used the following prompt so that the model is able to use chain-of-thought reasoning: "You can reason if you want but make sure to include yes/no in your answer." Interestingly, GPT4 performs much better than GPT3.5. We also observed that it is often the case that when GPT4 answers correctly to the question, it orders people based on their height, very similar to what we do in the scratchpad of the graph task. Motivated by this, we tested one more setting where we prompted GPT4 with "Answer only with a yes or no." to avoid the chain-of-thought reasoning. In this case, as expected, the model couldn’t solve the height comparison task for $n>1$. The results are shown in Figure 11.

### v2

@@ B.4 Length generalization for the parity task
where <pointer> determines the position of the $i$th bit and <value of the $i$th bit> is the value of the $i$th bit (which can be retrieved using the pointer) and <parity of the first $i$ bits> is the cumulative parity of the first $i$ bits that can be computed using its value in the previous state (parity of the first $i-1$ bits) and the value of the $i$th bit. For the last state, state $n+1$, we use [$d_{\mathrm{emb}}$]_,<final parity><EOS> where final parity is equal to the parity of the first $n$ bits that is computed in state $n$. So, one can easily check that this scratchpad is of low globality and each state can be easily computed using the previous state and question. As an example consider input _01_10_0__1_ again. The inductive scratchpad for this example can be written as
@@ B.5 Length generalization for the addition task
For both addition methods, we sampled numbers with $n\leq N_{\mathrm{train}}$ digits with a probability proportional to $n$. Nevertheless, we did not observe much dependency on the distribution. Also, note that in the design of scratchpads for both addition and parity, we have used techniques such as inserting random spaces and embedding in a fixed dimension. We note that this is unavoidable for the input as we have used absolute positional embedding for it. Nevertheless, we expect that by using an inductive scratchpad along with relative positional embedding and pre-training, one would be able to achieve the same length generalization with fewer assumptions on the format of the input.
@@ C.1 Architecture and datasets
For the cycle task, we use $1000$ node names formatted like v123. For simplicity of analysis, we regard each node name as a single token. Other than that and for the other tasks, we treat each character as a single token.
@@ C.3 Comparison with relative positional embeddings
Note that in the analysis above we assumed a fixed number of tokens per state. If the number of tokens in each state varies, issues like (2) and (3) above can become more difficult to handle for a model that only relies on relative positional embedding. To sum up, there is a minor similarity between relative positional embeddings and the idea of induction in the scratchpad. Factors such as pre-training and the recency bias of attention may promote inductive behavior in the model to some extent. Nevertheless, there is no reason to think that the model can implement the inductive behavior relying only on relative positional embeddings for a general inductive task. In general, one can use the inductive scratchpad idea along with relative positional embedding to get better length generalization. Note that inductive scratchpad can generalize on the number of training steps. However, understanding an input with growing length still requires relying on relative positional embeddings.
@@ Appendix E Comment on Lemma 1
Therefore, even for sets of size $n$, the mutual information is exponentially low, implying that $\mathrm{glob}(D)$ is greater than $n+1$.
@@ Appendix F Discussion on circuit complexity connections
Connection between globality degree and $NC^{0}$. We next show that, putting aside the histogram knowledge and using constant alphabet size, having low globality means having correlations with $NC^{0}$ circuits, a class of circuit weaker than $TC^{0}$ by constraining the number of fan-in to be constant (and not allowing for threshold gates).
@@
Let $P_{X}$ be a probability distribution over $\{0,1\}^{n}$ and $f\colon\{0,1\}^{n}\rightarrow\{0,1\}$ be a function. $f$ correlates non-trivially with a function computable in $NC^{0}$ if and only if $\mathrm{glob}(f)=O_{n}(1)$ when $X\sim P_{X}$.
@@
The $NC^{0}$ functions are exactly the binary functions that only depend on a constant number of input bits. So, for any $NC^{0}$ function $g$, there exists a set of input bits $S$ such that $|S|=O(1)$ and $g(X)$ only depends on the restriction of $X$ to $S$. That means that if $f$ has a nontrivial correlation with $g$ then $f(X)$ has a nontrivial correlation with $X[S]$ and thus has globality at most $|S|=O(1)$.
Now, assume that $\mathrm{glob}(f)=O_{n}(1)$ and choose $S$ with $|S|=O(1)$ such that $I(X[S],f(X))=n^{-O(1)}$. Next, define the function $g^{\prime}\colon\{0,1\}^{n}\rightarrow[0,1]$ such that $g^{\prime}(x)=\mathbb{E}[f(X)|X[S]=x[S]]$ for all $x$ and $g\colon\{0,1\}^{n}\rightarrow\{0,1\}$ such that
@@ Appendix G More experiments with ChatGPT
where the answer is true. Note that to answer this question correctly one has to combine at least $n+1$ relations. Thus, the globality of the task is always larger than $n$. (The exact globality would depend on the tokenization.) We found out that ChatGPT (GPT3.5) fails at this task even for $n=1$ (simplest case). Note that when working with the GPT3.5 model we used the following prompt so that the model is able to use chain-of-thought reasoning: "You can reason if you want but make sure to include yes/no in your answer." Interestingly, GPT4 performs much better than GPT3.5. We also observed that it is often the case that when GPT4 answers correctly to the question, it orders people based on their height, very similar to what we do in the scratchpad of the graph task. Motivated by this, we tested one more setting where we prompted GPT4 with "Answer only with a yes or no." to avoid the chain-of-thought reasoning. In this case, as expected, the model couldn’t solve the height comparison task for $n>1$. The results are shown in Figure 11.